Refer to the accompanying​ table, which describes results from groups of 8 births from 8 different sets of parents. The random variable x represents the number of girls among 8 children. Find the mean and standard deviation for the number of girls in 8 births

Number of Girls x P(x)
0 0.006
1 0.026
2 0.107
3 0.214
4 0.277
5 0.225
6 0.115
7 0.027
8 0.003


The mean is μ = ___ girls? Round to one decimal place as needed.
Answer #1
Occurance(X) Frequency(f) Freq*X (X-mean) (X-mean)2 f*(X-mean)2
0 0.006 0 -4.018 16.144 0.097
1 0.026 0.026 -3.018 9.108 0.237
2 0.107 0.214 -2.018 4.072 0.436
3 0.214 0.642 -1.018 1.036 0.222
4 0.277 1.108 -0.018 0 0
5 0.225 1.125 0.982 0.964 0.217
6 0.115 0.69 1.982 3.928 0.452
7 0.027 0.189 2.982 8.892 0.24
8 0.003 0.024 3.982 15.856 0.048
Total -> 1 4.018 - -

1.948

mean =EX =\sum x*p(x) = 4.018  ...(ANSWER)

Var = (\sum X^2*P(x))-(EX)^2

= 1.948

SD=\sqrt{variance} = \sqrt{1.948} = 1.396 ......(ANSWER